To go from your house to a nearby hospital, you must walk 11 m E and then turn 13 m 63º S of W. What is your displacement? Use graphical (parallelogram method) and trigonometric method.


1
Expert's answer
2021-10-03T13:24:02-0400

Trigonometric:


"D_E=11-13\cos63°=5.1\text m.\\
D_N=0-13\sin63°=-12\text m.\\
D=\sqrt{D_E^2+D_N^2}=13\text m.\\\space\\
\theta=\arctan\frac{D_N}{D_E}=70°\text{ S of E}."